# WVSC: Wireless Video Semantic Communication with Multi-frame Compensation

### v1

@@ IV-A4 Evaluation Metrics
(a) PNSR for the reconstructed videos.
(b) MS-SSIM for the reconstructed videos.
(c) LPIPS for the reconstructed videos.
@@ IV-B1 Performance for Different SNRs
We first testify the anti-noise performances of WVSC under Rayleigh fading channels with a specific channel bandwidth ratio (CBR), where CBR refers to the bandwidth compression ratio between the transmitted semantic codewords and the original video signals. We set it as 0.10. As shown in Fig. 5(a), it is clearly to observe that WVSC outperforms all other benchmarks with both WVSC-equal and WVSC-protected. Here WVSC-equal and WVSC-protected refer to the 1:1 and 4:1 rate allocation ratio between the reference frame and each residual, respectively. Compared to other DL-based schemes, WVSC outperforms DVSC for about 1 dB in terms of PSNR. This trend verifies that semantic-level wireless video transmission structure seems to be more robust to the channel fading and noise. For the traditional separated coding schemes, we adapt the ‘H.265+LDPC2/3+4QAM’ and the ‘H.264+LDPC2/3+4QAM’ as benchmarks, where LDPC code with 2/3 rate and QAM with modulation order 4 are set. Compared to traditional schemes, WVSC provides much more performance gain and stability since traditional schemes would be confronted with serious cliff effect with harsh channel conditions. As shown in Fig. 5(b) and Fig. 5(c), the DL-based schemes achieves better reconstruction results in terms of MS-SSIM and LPIPS compared to traditional separated coding schemes, which means that the satisfying visual perception quality is ensured through extracting semantics inside the original video signals. Compared to DVSC, WVSC preserves even much more high frequency image details. These results demonstrate the effectiveness of WVSC for video transmission under Rayleigh fading channels with various noise intensities.
@@ IV-B2 Performance for Different CBRs
Then we evaluate the bandwidth compression performances of WVSC under Rayleigh fading channels with SNR = 12 dB. As shown in Fig. 6(a), WVSC achieves much performance gain compared to other schemes. Compared to the DVSC, the performance gap increases as the increase of CBRs. Such insight mainly lies in the deployment of video coding in semantic level. Since semantic coding has great potential of source compression, based on the extracted semantics, the video coding enables much more effective data compression performances compared to directly conduct on video signals. In this way, higher CBRs allow the video coding part to retain more semantic information in the form of frame residuals. As the reference frames stay relatively steady as CBR values change, the increased allocation of extra bandwidth resources for transmitting residuals greatly promotes the reconstruction of each semantic frame. For the visual perceptual-level indexes in Fig. 6(b) and Fig. 6(c), WVSC retains relatively satisfying visual quality for a wide range of CBRs compared to other schemes.
Figure 7: Visualized results for the WVSC and other benchmarks.
IV-B3 Visualization Results for the Wireless Video Transmission
We present the visualization results for the wireless video transmission in Fig 7, where images are randomly picked from the reconstructed video clips. We select two types of separated coding schemes with different LDPC coding rates and modulation orders. In general, DL-based schemes outperform the traditional schemes for not only evaluation indexes such as PSNR and MS-SSIM but also visual quality since neural networks can retain more high frequency video signal details. It is evidently to observe that traditional schemes are unable to avoid blurry points for reconstructed videos while DL-based schemes provide more smooth images for different video clips. With proposed WVSC, reconstructed videos with sound visual reconstructed quality are provided.
@@ IV-B3 Visualization Results for the Wireless Video Transmission
IV-B4 Sensitive Analysis of different GoP Sizes

### v2

@@ IV-B1 Performance for Different SNRs
We first evaluate the anti-noise performances of WVSC under Rayleigh fading channels with a specific channel bandwidth ratio (CBR), where $CBR=\frac{L+(N-1)\times L_{1}}{N\times H\times W\times 3}$. We set it as 0.04. As shown in Fig. 5(a), it is clearly to observe that WVSC outperforms all other benchmarks with both WVSC and WVSC-protected. Here WVSC and WVSC-protected refer to the 1:1 and 4:1 rate allocation ratio between the I frame and each P frame, respectively. Compared to other DL-based schemes, WVSC outperforms DVSC for about 1 dB in terms of PSNR. This trend verifies that semantic-level wireless video transmission structure seems to be more robust to the channel fading and noise. For the traditional separated coding schemes, two channel coding and modulation parameter pairs are presented. In detail, 2/3 code rate LDPC with 4QAM and 1/2 code rate LDPC with 16QAM are set, which reflect corresponding anti-noise performances under different channel interference levels. Compared to traditional schemes, WVSC provides much more performance gain and stability since traditional schemes would be confronted with serious cliff effect with harsh channel conditions. As shown in Fig. 5(b) and Fig. 5(c), the DL-based schemes achieves better reconstruction results in terms of MS-SSIM and LPIPS compared to traditional SSCC schemes, which means that the satisfying visual perception quality is ensured through extracting semantics inside the original video signals. Compared to DVSC, WVSC preserves even much more high frequency image details. These results demonstrate the effectiveness of WVSC for video transmission under Rayleigh fading channels with various noise intensities.
@@ IV-B2 Performance for Different CBRs
Then we evaluate the bandwidth compression performances of WVSC under Rayleigh fading channels with SNR = 12 dB. As shown in Fig. 6(a), WVSC achieves significant performance gain compared to other schemes. Compared to the DVSC, the performance gap increases as CBR increases. Such insight mainly lies in the deployment of video coding in semantic level. Since semantic coding has great potential of source compression, based on the extracted semantics, the video coding enables much more effective data compression performances compared to directly conduct on video signals. In this way, higher CBRs allow the video coding part to retain more semantic information in the form of frame residuals. As the reference frames stay relatively steady as CBR values change, the increased allocation of extra bandwidth resources for transmitting residuals greatly promotes the reconstruction of each semantic frame. For the visual perceptual-level indexes in Fig. 6(b) and Fig. 6(c), WVSC retains relatively satisfying visual quality for a wide range of CBRs compared to other schemes.
@@
IV-B3 Sensitive Analysis of different GoP Sizes